OpenStudy (anonymous):

What is the general equation of parabola whose axis of symmetry is parallel to y-axis? ► ► ► ►

OpenStudy (anonymous):

your question is nonsensical

OpenStudy (anonymous):

y=ax^2 +b

OpenStudy (anonymous):

y=ax^2 +bx+c which one

OpenStudy (unklerhaukus):

y=ax^2 +b

OpenStudy (mathmate):

A parabola can have the axis parallel to the x-axis, even though the complete parabola does not make a function (because it does not pass the vertical line test). The general formula for the parabola with a vertical axis is: y=a(x-h)+k where (h,k) is the vertex. so the one with a horizontal axis is x=a(y-k)+h where (h,k) is still the vertex.